### v4

@@ Abstract
  • MGS for groups belongs to $\mathsf{AC}^{1}(\mathsf{L})$. Our $\mathsf{AC}^{1}(\mathsf{L})$ bound improves on the previous, very recent, upper bound of P (Lucchini & Thakkar, J. Algebra, 2024). Our $\mathsf{quasiAC^{0}}$ upper bound is incomparable to $\mathsf{P}$, but has similar consequences to the above result for quasigroups.
  • Quasigroup Isomorphism belongs to $\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))\subseteq\mathsf{quasiAC^{0}}$. As a consequence of this result and previously known $\mathsf{AC^{0}}$ reductions, this implies the same upper bound for the Isomorphism Problems for: Steiner triple systems, pseudo-STS graphs, Latin Square Isotopy, Latin square graphs, and Steiner $(t,t+1)$-designs. This improves upon the previous upper bound for these problems, which was $\exists^{\log^{2}n}\mathsf{L}\cap\exists^{\log^{2}n}\mathsf{FOLL}\subseteq\mathsf{quasiFOLL}$ (Chattopadhyay, Torán, & Wagner, ibid.; Levet, Australas. J. Combin. 2023).
@@ 1 Introduction
The Group Isomorphism (GpI) problem is a central problem in computational complexity and computer algebra. When the groups are given as input by their multiplication (a.k.a. Cayley) tables, the problem reduces to Graph Isomorphism (GI), and because the best-known runtimes for the two are quite close ($n^{O(\log n)}$ [Mil78] [^2] Miller [Mil78] credits Tarjan for $n^{\log n+O(1)}$. vs. $n^{O(\log^{2}n)}$ [Bab16] [^3] Babai [Bab16] proved quasi-polynomial time, and the exponent of the exponent was analyzed and improved by Helfgott [HBD17]), the former stands as a key bottleneck towards further improvements in the latter.
Despite this, GpI seems quite a bit easier than GI. For example, Tarjan’s $n^{\log n+O(1)}$ algorithm for groups [Mil78] can now be given as an exercise to undergraduates: every group is generated by at most $\lceil\log_{2}|G|\rceil$ elements, so the algorithm is to try all possible $\binom{n}{\log n}\leq n^{\log n}$ generating sets, and for each, check in $n^{O(1)}$ time whether the map of generating sets extends to an isomorphism. In contrast, the quasi-polynomial time algorithm for graphs was a tour de force that built on decades of research into algorithms and the structure of permutation groups. Nonetheless, it remains unknown whether the problem for groups is actually easier than that for graphs under polynomial-time reductions, or even whether both problems are in $\mathsf{P}$!
Using a finer notion of reduction, Chattopadhyay, Torán, and Wagner [CTW13] proved that there was no $\mathsf{AC^{0}}$ reduction from GI to GpI. This gave the first unconditional evidence that there is some formal sense (namely, the $\mathsf{AC^{0}}$ sense) in which GpI really is easier than GI. The key to their result was that the generator-enumeration technique described above can be implemented by non-deterministically guessing $\log^{2}n$ bits (describing the $\log n$ generators, each of $\log n$ bits), and then verifying an isomorphism by a non-deterministic circuit of depth only $O(\log\log n)$, a class we denote $\exists^{\log^{2}n}\mathsf{FOLL}$. Observe that $\exists^{\log^{2}n}\mathsf{FOLL}\subseteq\textsf{quasiFOLL}$ (by trying all $2^{O(\log^{2}n)}$ settings of the non-deterministic bits in parallel), which cannot compute Parity [Hås86], even if augmented with $MOD_{p}$ gates for $p$ an odd prime [Raz93, Smo87, CTW13]. As GI is $\mathsf{DET}$-hard [Tor04]—and hence can compute Parity—there can be no $\mathsf{AC^{0}}$ reduction from GI to GpI.
@@
The authors would be shocked if the answer were “yes,” and yet we do not even have results showing that Group Isomorphism cannot be computed by polynomial-size circuits of (!) depth 2. Indeed, the upper bound of $\exists^{\log^{2}n}\mathsf{FOLL}$ rules out most existing lower bound techniques against $\mathsf{AC^{0}}$, as most such techniques also yield similar lower bounds against $\exists^{\log^{2}n}\mathsf{FOLL}$.
In this paper, we aim to close the gap between $\mathsf{AC^{0}}$ and $\exists^{\log^{2}n}\mathsf{FOLL}$ in the complexity of Group Isomorphism and related problems. Our goal is to obtain constant-depth circuits of quasipolynomial size, a natural benchmark in circuit complexity [Bar92]. Significantly improving the size of these circuits would improve the state of the art run-time of Group Isomorphism, a long-standing open question that we do not address here.
@@
We in fact get a more precise bound of $\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$ (Theorem 6.1 gives an even more specific bound). This more precise bound is notable because it is contained in both $\mathsf{quasiAC^{0}}$ and $\exists^{\log^{2}n}\mathsf{FOLL}\cap\exists^{\log^{2}n}\mathsf{L}$, the latter thus improving on [CTW13]. We get similarly precise bounds with complicated-looking complexity classes for the other problems mentioned in the introduction, but we omit the precise bounds here for readability.
Focusing on depth bounds—that is, without attempting to improve the worst-case runtime—our result is close to the end of the line for a series of works stretching back to 1970; see Table 1. The only possible further improvements we see, without improving the worst-case runtime, are to get an $\exists^{\log^{2}n}\mathsf{AC^{0}}$ upper bound—for which there are many obstacles, see Section 8—or improving the exact size and depth of our result, but here there is not much room for improvement, as we already get quasi-polynomial-size circuits of depth only 4 and size $n^{O(\log^{c}n)}$ for some $1<c\leq 4$.
@@
1970 Generator-enumerator introduced $\operatorname{poly}(n)$ Felsch & Neubüser [FN70] [^4] Complexity not analyzed there, but the same as Tarjan’s algorithm [Mil78].
1978 $\exists^{\log^{2}n}\mathsf{P}\subseteq\mathsf{DTIME}(n^{\log n+O(1)})$ $\operatorname{poly}(n)$ Tarjan (see Miller [Mil78])
1977 $\mathsf{DSPACE}(\log^{2}n)$ $O(\log^{2}n)$ Lipton–Snyder–Zalcstein [LSZ77] [^5] Despite the publication dates, this seems to have been independent of Tarjan’s result. They note that Miller and Rabin had also observed this result independently.
1994 $\exists^{\log^{2}n}\mathsf{AC}^{1}$ $O(\log n)$ Wolf [Wol94] [^6] Wolf only claims a bound of $\exists^{\log^{2}n}\textsf{NC}^{2}$. However, if we replace his use of $\textsf{NC}^{1}$ circuits to multiply two elements of a quasigroup with $\mathsf{AC^{0}}$ circuits, we immediately get the $\exists^{\log^{2}n}\mathsf{AC}^{1}$ bound.
2010 $\exists^{\log^{2}n}\mathsf{SAC}^{1}$ $O(\log n)$ Wagner [Wag10]
2010 $\exists^{\log^{2}n}\mathsf{FOLL}\cap\exists^{\log^{2}n}\mathsf{L}$ $O(\log\log n)$ Chattopadhyay–Torán–Wagner [CTW13] [^7] They do not claim the $\exists^{\log^{2}n}\mathsf{L}$ bound, but it follows immediately from their algorithm and results.
2013 $\exists^{\log^{2}n}\mathsf{SC}^{2}\cap\exists^{\log^{2}n}\mathsf{L}$ $O(\log n)$ Papakonstantinou–Tang–Qiao [Tan13] [^8] We have written the result this way, despite $\exists^{\log^{2}n}\mathsf{L}\subseteq\exists^{\log^{2}n}\mathsf{SC}^{2}$, because in Tang’s thesis [Tan13], the only place this is currently published, they only claim $\mathsf{NSC}^{2}$ using only $O(\log^{2}n)$ bits of nondeterminism, which in our notation would be $\exists^{\log^{2}n}\mathsf{SC}^{2}$. However, their algorithm and results also immediately yields a $\exists^{\log^{2}n}\mathsf{L}$ bound.
@@
Table 1: History of the low-level circuit complexity of algorithms for Group Isomorphism based on the generator-enumerator technique. For non-circuit classes, we list their depth as the best-known depth of their simulation by circuits. The class in our bound is contained in all the other classes listed in the table. Although depth 4 in our result does not follow from the complexity class as listed here, it follows from the more exact class we use in Theorem 6.1.
We note that although there are depth reduction techniques for bounded-depth circuits with mod-$p$ gates [AH94], no such result is known for $\mathsf{quasiAC^{0}}$ circuits (without mod-$p$ gates).
@@
Another very natural problem in computational algebra is the Min Generating Set (MGS) problem. Given a group, this problem asks to find a generating set of the smallest possible size. Given that many algorithms on groups depend on the size of a generating set, finding a minimum generating set has the potential to be a widely applicable subroutine. The MGS problem for groups was shown to be in $\mathsf{P}$ by Lucchini & Thakkar only very recently [LT24]. We improve their complexity bound to:
Theorem B.
@@
While our $\mathsf{AC}^{1}(\mathsf{L})$ bound above is essentially a careful complexity analysis of the polynomial-time algorithm of Lucchini & Thakkar [LT24], the $\mathsf{quasiAC^{0}}$ upper bound is in fact a consequence of our next, more general result for quasigroups, which involves some new ingredients.
@@
Quasigroups can be defined in (at least) two equivalent ways: (1) an algebra whose multiplication table is a Latin square, [^9] A Latin square is an $n\times n$ matrix where for each row and each column, the elements of $[n]$ appear exactly once or (2) a group-like algebra that need not have an identity nor be associative, but in which left and right division are uniquely defined, that is, for all $a,b$, there are unique $x$ and $y$ such that $ax=b$ and $ya=b$.
In the paper in which they introduced $\log^{2}(n)$-bounded nondeterminism, Papadimitriou and Yannakakis showed that for arbitrary magmas [^10] A magma is a set $M$ together with a function $M\times M\to M$ that need not satisfy any additional axioms., testing whether the magma has $\log n$ generators was in fact complete for $\exists^{\log^{2}n}\mathsf{P}$, and conjectured:
Conjecture (Papadimitriou & Yannakakis [PY96, p. 169]).
@@
“We conjecture that this result [$\exists^{\log^{2}n}\mathsf{P}$-completeness] also holds for the more structured MINIMUM GENERATOR SET OF A QUASIGROUP problem. In contrast, QUASIGROUP ISOMORPHISM was recently shown to be in DSPACE$(\log^{2}n)$ [Wol94]. Notice that the corresponding problems for groups were known to be in DSPACE$(\log^{2}n)$ [LSZ77].”—Papadimitriou & Yannakakis [PY96, p. 169]
We thus turn our attention to the analogous problems for quasigroups: MGS for quasigroups, Quasigroup Isomorphism, and the key subroutine, Sub-quasigroup Membership. We note that the $\exists^{\log^{2}n}\mathsf{FOLL}$ upper bound of Chattopadhyay, Torán, and Wagner [CTW13] actually applies to Quasigroup Isomorphism and not just GpI; we perform a careful analysis of their algorithm to put Quasigroup Isomorphism into $\mathsf{quasiAC^{0}}$ as well.
Theorem C.
@@
To the best of our knowledge, MGS for Quasigroups has not been studied from the complexity-theoretic viewpoint previously. While a $\mathsf{DSPACE}(\log^{2}n)$ upper bound for MGS for groups follows from [Tan13, AT06], as far as we know it remained open for quasigroups prior to our work.
As with prior results on Quasigroup Isomorphism and Group Isomorphism [CTW13], and other isomorphism problems such as Latin Square Isotopy and Latin Square Graph Isomorphism [Lev23], Thm. C shows that Parity does not reduce to MGS for Quasigroups, thus ruling out most known lower bound methods that might be used to prove that MGS for Quasigroups is not in $\mathsf{AC^{0}}$. We also observe a similar bound for MGS for Groups using Fleischer’s technique [Fle22].
@@
Corollary C.
The conjecture of Papadimitriou & Yannakakis [PY96, p. 169] is false under $\mathsf{quasiAC^{0}}$ reductions. It is also false under polylog-space reductions assuming $\mathsf{EXP}\neq\mathsf{PSPACE}$.
In strong contrast, we show that MGS for Magmas is $\mathsf{NP}$-complete (Thm. 7.12).
A key ingredient in our proof of Thm. C is an improvement in the complexity of another central problem in computational algebra: the Sub-quasigroup Membership problem (Membership, [^11] In the literature, the analogous problem for groups is sometimes called Cayley Group Membership or CGM, to highlight that it is in the Cayley table model. for short):
Theorem D.
@@
Membership for groups is well-known to belong to $\mathsf{L}$, by reducing to the connectivity problem on the appropriate Cayley graph (cf. [BM91, Rei08]), but as $\mathsf{L}$ sits in between $\mathsf{AC^{0}}$ and $\mathsf{AC}^{1}$, this is not low enough depth for us.
@@
  • By known $\mathsf{AC^{0}}$ reductions (see, e.g., Levet [Lev23] for details), our $\mathsf{quasiAC^{0}}$ analysis of Chattopadhyay, Torán, and Wagner’s algorithm for Quasigroup Isomorphism yields the same upper bound for the isomorphism problems for Steiner triple systems, pseudo-STS graphs, Latin square graphs, and Steiner $(t,t+1)$-designs, as well as Latin Square Isotopy. Prior to our work, Quasigroup Isomorphism was not known to be solvable using quasiAC circuits of depth $o(\log\log n)$. See Cor. 6.2.
  • Group Isomorphism for simple groups (Cor. 3.3) or for groups from a dense set $\Upsilon$ of orders (Thm. 5.1) can be solved in $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))\subseteq\mathsf{quasiAC^{0}}\cap\mathsf{L}\cap\mathsf{FOLL}$. For groups in a dense set of orders, this improves the parallel complexity compared to the original result of Dietrich & Wilson [DW22]. As in their paper, note that $\Upsilon$ omits large prime powers. Thus, we essentially have that for groups that are not $p$-groups, Group Isomorphism belongs to a proper subclass of DET. This evidence fits with the widely-believed idea that $p$-groups are a bottleneck case for Group Isomorphism.
  • Abelian Group Isomorphism (Thm. 4.1) is in $\forall^{\log\log n}\mathsf{MAC}^{0}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$. The key novelties here are (1) a new observation that allows us to reduce the number of co-nondeterministic bits from $\log n$ (as in [GL23]) down to $\log\log n$, and (2) using an $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$ circuit for order finding, rather than $\mathsf{FOLL}$ as in [CTW13].
  • Membership for nilpotent groups is in $\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n))\subseteq\mathsf{FOLL}\cap\mathsf{quasiAC^{0}}$ (Prop. 7.3).

### v5

@@ Abstract
  • MGS for groups belongs to $\mathsf{AC}^{1}(\mathsf{L})$. Our $\mathsf{AC}^{1}(\mathsf{L})$ bound improves on the previous, very recent, upper bound of $\mathsf{P}$ (Lucchini & Thakkar, J. Algebra, 2024). Our $\mathsf{quasiAC^{0}}$ upper bound is incomparable to $\mathsf{P}$, but has similar consequences to the above result for quasigroups.
  • Quasigroup Isomorphism belongs to $\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathsf{TIMESPACE}(\text{polylog}(n),\log(n)))\subseteq\mathsf{quasiAC^{0}}$. As a consequence of this result and previously known $\mathsf{AC^{0}}$ reductions, this implies the same upper bound for the Isomorphism Problems for: Steiner triple systems, pseudo-STS graphs, Latin Square Isotopy, Latin square graphs, and Steiner $(t,t+1)$-designs. This improves upon the previous upper bound for these problems, which was $\exists^{\log^{2}n}\mathsf{L}\cap\exists^{\log^{2}n}\mathsf{FOLL}\subseteq\mathsf{quasiFOLL}$ (Chattopadhyay, Torán, & Wagner, ibid.; Levet, Australas. J. Combin. 2023).
@@ 1 Introduction
The Group Isomorphism (GpI) problem is a central problem in computational complexity and computer algebra. When the groups are given as input by their multiplication (a.k.a. Cayley) tables, the problem reduces to Graph Isomorphism (GI), and because the best-known runtimes for the two are quite close ($n^{O(\log n)}$ [43] [^1] Miller [43] credits Tarjan for $n^{\log n+O(1)}$. vs. $n^{O(\log^{2}n)}$ [5] [^2] Babai [5] proved quasi-polynomial time, and the exponent of the exponent was analyzed and improved by Helfgott [29].), the former stands as a key bottleneck towards further improvements in the latter.
Despite this, GpI seems quite a bit easier than GI. For example, Tarjan’s $n^{\log n+O(1)}$ algorithm for groups [43] can now be given as an exercise to undergraduates: every group is generated by at most $\lfloor\log_{2}|G|\rfloor$ elements, so the algorithm is to try all possible $\binom{n}{\log n}\leq n^{\log n}$ generating sets, and for each, check in $n^{O(1)}$ time whether the map of generating sets extends to an isomorphism. In contrast, the quasi-polynomial time algorithm for graphs was a tour de force that built on decades of research into algorithms and the structure of permutation groups. Nonetheless, it remains unknown whether the problem for groups is actually easier than that for graphs under polynomial-time reductions, or even whether both problems are in $\mathsf{P}$!
Using a finer notion of reduction, Chattopadhyay, Torán, and Wagner [17] proved that there was no $\mathsf{AC^{0}}$ reduction from GI to GpI. This gave the first unconditional evidence that there is some formal sense (namely, the $\mathsf{AC^{0}}$ sense) in which GpI really is easier than GI. The key to their result was that the generator-enumeration technique described above can be implemented by non-deterministically guessing $\log^{2}n$ bits (describing the $\log n$ generators, each of $\log n$ bits), and then verifying an isomorphism by a non-deterministic circuit of depth only $O(\log\log n)$, a class we denote $\exists^{\log^{2}n}\mathsf{FOLL}$. Observe that $\exists^{\log^{2}n}\mathsf{FOLL}\subseteq\mathsf{quasiFOLL}$ (by trying all $2^{O(\log^{2}n)}$ settings of the non-deterministic bits in parallel), which cannot compute Parity [28], even if augmented with $\mathsf{Mod}_{p}$ gates for $p$ an odd prime [46, 52, 17]. As GI is $\mathsf{DET}$-hard [55]—and hence can compute Parity—there can be no $\mathsf{AC^{0}}$ reduction from GI to GpI.
@@
The authors would be shocked if the answer were “yes,” and yet we do not even have results showing that Group Isomorphism cannot be computed by polynomial-size circuits of (!) depth 2. Indeed, it is not clear how to use existing $\mathsf{AC^{0}}$ lower bound techniques against Group Isomorphism. [^3] The $\exists^{\log^{2}n}\mathsf{FOLL}$ upper bound unconditionally rules out reductions from Parity and Majority. While switching lemmas have been used to get $\mathsf{AC^{0}}$ lower bounds on LogClique (deciding if a graph has a clique on $O(\log n)$ vertices) [42, 14, 50] (covered in Beame’s switching lemma primer [13]), which is in $\exists^{\log^{2}n}\mathsf{AC^{0}}\subseteq\mathsf{quasiAC^{0}}$, that problem feels quite different from Group Isomorphism.
In this paper, we aim to close the gap between $\mathsf{AC^{0}}$ and $\exists^{\log^{2}n}\mathsf{FOLL}$ in the complexity of Group Isomorphism and related problems. Our goal is to obtain constant-depth circuits of quasipolynomial size, a natural benchmark in circuit complexity [9]. Significantly improving the size of these circuits would improve the state of the art run-time of Group Isomorphism, a long-standing open question that we do not address here.
@@
We in fact get a more precise bound of $\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$ (Theorem 6.1 gives an even more specific bound), where $\mathsf{DTISP}(t(n),s(n))$ is the class of languages decidable by a Turing machine that simultaneously uses time at most $t$ and space at most $s$. This more precise bound is notable because it is contained in both $\mathsf{quasiAC^{0}}$ and $\exists^{\log^{2}n}\mathsf{FOLL}\cap\exists^{\log^{2}n}\mathsf{L}$, the latter thus improving on [17]. We get similarly precise bounds with complicated-looking complexity classes for the other problems mentioned in the introduction, but we omit the precise bounds here for readability.
Focusing on depth bounds—that is, without attempting to improve the worst-case runtime—our result is close to the end of the line for a series of works stretching back to 1970; see Table 1. The only possible further improvements we see, without improving the worst-case runtime, are to get an $\exists^{\log^{2}n}\mathsf{AC^{0}}$ upper bound—for which there are several obstacles, see Section 8—or improving the exact size and depth of our result, but there is not much room for improvement here, as we already get quasi-polynomial-size circuits of depth only 4 and size $n^{O(\log n)}$, matching the current-best serial runtime up to the constant in the exponent (we have not attempted to optimize the constant hidden in the big-Oh; we certainly do not get a constant less than 1, and we conservatively estimate our proof yields a constant not more than $20$).
@@
1970 Generator-enumerator introduced $\operatorname{poly}(n)$ Felsch & Neubüser [23] [^4] Complexity not analyzed there, but the same as Tarjan’s algorithm [43].
1978 $\exists^{\log^{2}n}\mathsf{P}\subseteq\mathsf{DTIME}(n^{\log n+O(1)})$ $\operatorname{poly}(n)$ Tarjan (see Miller [43])
1977 $\mathsf{DSPACE}(\log^{2}n)$ $O(\log^{2}n)$ Lipton–Snyder–Zalcstein [38] [^5] Despite the publication dates, this seems to have been independent of Tarjan’s result. They note that Miller and Rabin had also observed this result independently.
1994 $\exists^{\log^{2}n}\mathsf{AC}^{1}$ $O(\log n)$ Wolf [59] [^6] Wolf only claims a bound of $\exists^{\log^{2}n}\mathsf{NC}^{2}$. However, if we replace his use of $\mathsf{NC}^{1}$ circuits to multiply two elements of a quasigroup with $\mathsf{AC^{0}}$ circuits, we immediately get the $\exists^{\log^{2}n}\mathsf{AC}^{1}$ bound.
2010 $\exists^{\log^{2}n}\mathsf{SAC}^{1}$ $O(\log n)$ Wagner [58]
2010 $\exists^{\log^{2}n}\mathsf{FOLL}\cap\exists^{\log^{2}n}\mathsf{L}$ $O(\log\log n)$ Chattopadhyay–Torán–Wagner [17] [^7] They do not claim the $\exists^{\log^{2}n}\mathsf{L}$ bound, but it follows immediately from their algorithm and results.
2013 $\exists^{\log^{2}n}\mathsf{SC}^{2}\cap\exists^{\log^{2}n}\mathsf{L}$ $O(\log n)$ Papakonstantinou–Tang–Qiao [53] [^8] We have written the result this way, despite $\exists^{\log^{2}n}\mathsf{L}\subseteq\exists^{\log^{2}n}\mathsf{SC}^{2}$, because in Tang’s thesis [53], the only place this is currently published, they only claim $\mathsf{NSC}^{2}$ using only $O(\log^{2}n)$ bits of nondeterminism, which in our notation would be $\exists^{\log^{2}n}\mathsf{SC}^{2}$. However, their algorithm and results also immediately yields a $\exists^{\log^{2}n}\mathsf{L}$ bound.
@@
Table 1: History of the low-level circuit complexity of algorithms for (Quasi)Group Isomorphism based on the generator-enumerator technique. For non-circuit classes, we list their depth as the best-known depth of their simulation by circuits. The class in our bound is contained in all the other classes listed in the table. Although depth 4 in our result does not follow from the complexity class as listed here, it follows from the more exact class we use in Theorem 6.1.
We note that although there are depth reduction techniques for bounded-depth circuits with $\mathsf{Mod}_{p}$ gates [2], no such result is known for $\mathsf{quasiAC^{0}}$ circuits (without $\mathsf{Mod}_{p}$ gates).
@@
Another very natural problem in computational algebra is the Min Generating Set (MGS) problem. Given a group, this problem asks to find a generating set of the smallest possible size. Given that many algorithms on groups depend on the size of a generating set, finding a minimum generating set has the potential to be a widely applicable subroutine. The MGS problem for groups was shown to be in $\mathsf{P}$ by Lucchini & Thakkar only very recently [40]. We improve their complexity bound to:
Theorem 1.3.
@@
While our $\mathsf{AC}^{1}(\mathsf{L})$ bound above is essentially a careful complexity analysis of the polynomial-time algorithm of Lucchini & Thakkar [40], the $\mathsf{quasiAC^{0}}$ upper bound is in fact a consequence of our next, more general, result for quasigroups, which involves some new ingredients.
@@
Quasigroups can be defined in (at least) two equivalent ways: (1) an algebra whose multiplication table is a Latin square, [^9] A Latin square is an $n\times n$ matrix where for each row and each column, the elements of $[n]$ appear exactly once. or (2) a group-like algebra that need not have an identity nor be associative, but in which left and right division are uniquely defined, that is, for all $a,b$, there are unique $x$ and $y$ such that $ax=b$ and $ya=b$.
In the paper in which they introduced $\log^{2}(n)$-bounded nondeterminism, Papadimitriou and Yannakakis showed that for arbitrary magmas, [^10] A magma is a set $M$ together with a function $M\times M\to M$ that need not satisfy any additional axioms. testing whether the magma has $\log n$ generators was in fact complete for $\exists^{\log^{2}n}\mathsf{P}$, and conjectured:
Conjecture 1.4 (Papadimitriou & Yannakakis [45, p. 169]).
@@
[…]
We thus turn our attention to the analogous problems for quasigroups: MGS for quasigroups, Quasigroup Isomorphism, and the key subroutine, Sub-quasigroup Membership. We note that the $\exists^{\log^{2}n}\mathsf{FOLL}$ upper bound of Chattopadhyay, Torán, and Wagner [17] actually applies to Quasigroup Isomorphism and not just GpI; we perform a careful analysis of their algorithm to put Quasigroup Isomorphism into $\mathsf{quasiAC^{0}}$ as well.
Theorem 1.5.
@@
To the best of our knowledge, MGS for Quasigroups has not been studied from the complexity-theoretic viewpoint previously. While a $\mathsf{DSPACE}(\log^{2}n)$ upper bound for MGS for groups follows from [53, 3], as far as we know it remained open for quasigroups prior to our work.
As with prior results on Quasigroup Isomorphism and Group Isomorphism [17], and other isomorphism problems such as Latin Square Isotopy and Latin Square Graph Isomorphism [37], Thm. 1.5 implies that Parity does not reduce to MGS for Quasigroups, thus ruling out most known lower bound methods that might be used to prove that MGS for Quasigroups is not in $\mathsf{AC^{0}}$. We also observe a similar bound for MGS for Groups using Fleischer’s technique [25].
@@
Corollary 1.6.
The conjecture of Papadimitriou & Yannakakis [45, p. 169] is false under $\mathsf{quasiAC^{0}}$ reductions. It is also false under polylog-space reductions assuming $\mathsf{EXP}\neq\mathsf{PSPACE}$.
In strong contrast, we show that MGS for Magmas is $\mathsf{NP}$-complete (Thm. 7.19).
A key ingredient in our proof of Thm. 1.5 is an improvement in the complexity of another central problem in computational algebra: the Sub-quasigroup Membership problem (Membership, [^11] In the literature, the analogous problem for groups is sometimes called Cayley Group Membership or CGM, to highlight that it is in the Cayley table model. for short):
Theorem 1.7.
@@
Membership for groups is well-known to belong to $\mathsf{L}$, by reducing to the connectivity problem on the appropriate Cayley graph (cf. [11, 47]), but as $\mathsf{L}$ sits in between $\mathsf{AC^{0}}$ and $\mathsf{AC}^{1}$, this is not low enough depth for us.
@@
  • By known $\mathsf{AC^{0}}$ reductions (see, e.g., Levet [37] for details), our $\mathsf{quasiAC^{0}}$ analysis of Chattopadhyay, Torán, and Wagner’s algorithm for Quasigroup Isomorphism yields $\mathsf{quasiAC^{0}}$ upper bound for the isomorphism problems for Steiner triple systems, pseudo-STS graphs, Latin square graphs, and Steiner $(t,t+1)$-designs, as well as Latin Square Isotopy. Prior to our work, Quasigroup Isomorphism was not known to be solvable using $\mathsf{quasiAC}$ circuits of depth $o(\log\log n)$. See Cor. 6.3.
  • Group Isomorphism for simple groups (Cor. 3.5) or for groups from a dense set $\Upsilon$ of orders (Thm. 5.1) can be solved in $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))\subseteq\mathsf{quasiAC^{0}}\cap\mathsf{L}\cap\mathsf{FOLL}$. For groups in a dense set of orders, this improves the parallel complexity compared to the original result of Dietrich & Wilson [22]. As in their paper, note that $\Upsilon$ omits large prime powers. Thus, we essentially have that for groups that are not $p$-groups, Group Isomorphism belongs to a proper subclass of $\mathsf{DET}$. This evidence fits with the widely-believed idea that $p$-groups are a bottleneck case for Group Isomorphism.
  • Abelian Group Isomorphism (Thm. 4.1) is in $\forall^{\log\log n}\mathsf{MAC}^{0}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$. The key novelties here are (1) a new observation that allows us to reduce the number of co-nondeterministic bits from $\log n$ (as in [26]) down to $\log\log n$, and (2) using an
    $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$ circuit for order finding, rather than $\mathsf{FOLL}$ as in [17].
  • Membership for nilpotent groups is in $\mathsf{AC^{0}}(\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))\subseteq\mathsf{FOLL}\cap\mathsf{quasiAC^{0}}$ (Prop. 7.5).